# 6D Rigid Body Localization and Velocity Estimation via Gaussian Belief Propagation

### v2

@@ Abstract
Index Terms:
Rigid body localization (RBL), Message Passing, 6D Localization, Attitude Estimation.
@@ I Introduction
Several effective strategies exists to estimate the location and attitude of objects, including computer vision-based techniques, which focus on feature extraction and posture estimation [15, 16], usually based on image/video signals; and inertial measurement unit (IMU)-based techniques, which leverage information from accelerometers, gyroscopes, and magnetometers [17, 18]. A problem of computer-vision approaches is, however, that they typically require high volumes of data and rely on high-complexity methods, which limits their application. In turn, IMU-based methods are often unreliable or inaccurate, requiring frequent sensor re-calibrations, not to mention the aid of external radio technology, which to some extent defeats the self-reliant idea behind the approach.
In contrast to the latter, the RBL concept considered here [9, 10, 11, 12, 13, 14] aims to generalize the traditional localization paradigm, and exploits range measurements between sensors on the rigid body and a set of anchor sensors at known positions [19, 20], to estimate the location, orientation (and possibly the shape) of rigid bodies. Earlier state-of-the-art (SotA) contributions on such RBL approach are, however, based on either algebraic methods leveraging multidimensional scaling (MDS) [21] or semidefinite relaxation (SDR) [22, 23], both of which have a cubic computational complexity at least. And although least squares methods [24, 25, 26] have been used to lower the complexity of the RBL problem, these techniques have in common the fact that they only take into account stationary scenarios, in the sense that they seek to estimate only the translation and the orientation ($i.e.$, the pose) of the rigid body, respectively defined as the distance and the rotation matrix of the rigid body relative to a given reference.
@@
The capitalizing feature of the method in [11] to handle both the stationary and moving RBL problems similarly is thanks to the linearized error model described in [29] and [30]. However, the least-squares formulation and two stage approach adopted can limit performance [31] compared to message-passing methods [32], which are well known to achieve low complexity while maintaining a good performance especially in problems with coupled variables, which can be solved not only via alternating methods [33], but also jointly via bilinear techniques [34, 35, 36].
In light of the above, and inspired by former message passing based localization approaches [37, 38, 39], we propose here a series of novel algorithms to solve the stationary and moving RBL problems described in the article. First, two methods for the location and velocity estimation of sensors in three-dimensional (3D) space are produced. In particular, the techniques, labeled Algorithms 1 and 3, respectively, are designed for the estimation of sensor positions from range measurements, as well as sensor velocities from Doppler measurements. Then, two bivariate linear Gaussian belief propagation (GaBP) algorithms are offered for the stationary and moving RBL problems. Those contributions, dubbed Algorithms 2 and 4, respectively, are capable of estimating the 3D rotation angles and translation distances, and respectively the 3D angular and translational velocities, utilizing range/Doppler measurements and the estimated sensor position and velocities from Algorithms 1 and 2 by leveraging bilinear GaBP over a small-angle modification of the linearized system models of [30] and [29], with the rigid body conformation constraint directly incorporated. The bivariate methods in Algorithms 4 and 3, respectively, are shown to outperform the SotA two-stage RBL methods in terms of parameter estimation performance, while retaining low computational complexity.
@@ II-A Rigid Body System Model
Figure 2: An illustration of a moving rigid body, whose location has a 3D rotation $\bm{Q}$ and a translation $\bm{t}$ from the reference frame [^3] The reference frame is generally not at the origin, but since a rigid body rotation is only relative to its previous orientation, setting the reference frame to the origin, $i.e.$, making $\bm{Q}^{(0)}\triangleq\mathbf{I}_{3\times 3}$, can be done without loss of generality (w.l.g.)., as determined by equation (11). The velocity $\bm{\dot{s}}_{n}$ of the $n$-th sensor as a function of the angular velocity $\bm{\omega}$ and translational velocity $\bm{\dot{t}}$, as determined by equation (12), is also illustrated.
@@
The latter system model for stationary rigid bodies can be extended to moving rigid bodies by introducing the angular velocity vector $\bm{\omega}\triangleq\left[\omega_{1},\omega_{2},\omega_{3}\right]^{\mathsf{T}}\in\mathbb{R}^{3\times 1}$, and the translational velocity vector $\bm{\dot{t}}\triangleq[\dot{t}_{x},\dot{t}_{y},\dot{t}_{z}]^{\mathsf{T}}\in\mathbb{R}^{3\times 1}$, such that the velocity of the $n$-th sensor can be expressed as
@@
where $[\cdot]^{\times}$ is the cross product operator for matrices [10], which maps the angular velocity vector $\bm{\omega}$ to a skew-symmetric matrix, given by
$[\bm{\omega}]^{\times}=\left[\begin{array}[]{ccc}0&-\omega_{3}&\omega_{2}\\ \omega_{3}&0&-\omega_{1}\\ -\omega_{2}&\omega_{1}&0\end{array}\right]\in\mathbb{R}^{3\times 3}.$ (13)
@@ II-B Sensor Range Measurement Model
where the second-order noise term $w_{m,n}^{2}$ is neglected.
@@
$\bm{y}_{n}\!$ $\triangleq\!\!\!\begin{bmatrix}\!\tilde{d}_{1,n}^{2}\!\!-\!\left\|\bm{a}_{1}\right\|^{2}_{2}\!\\ \!\vdots\\ \tilde{d}_{M,n}^{2}\!\!-\!\left\|\bm{a}_{M}\right\|^{2}_{2}\\ \end{bmatrix}\!\!\!=\!\!\!\underbrace{\begin{bmatrix}\!-2\bm{a}_{1}^{\mathsf{T}},&\!\!\!\!1\;\\ \!\vdots&\!\!\!\!\vdots~{}\\ \!-2\bm{a}_{M}^{\mathsf{T}},&\!\!\!\!1\;\\ \end{bmatrix}}_{\triangleq\,\bm{G}\,\in\,\mathbb{R}^{M\!\times\!4}}\!\!\!\!\!\!\!\overbrace{\!\begin{bmatrix}\bm{s}_{n}\\[4.30554pt] ||\bm{s}_{n}||_{2}^{2}\end{bmatrix}\!}^{~{}~{}~{}\triangleq\,\bm{x}_{n}\,\in\,\mathbb{R}^{4\!\times\!1}}\!\!\!\!\!\!+\!\!\!\!\!\!\underbrace{\!\begin{bmatrix}\!\xi_{1,n}\!\\ \!\vdots\!\\ \!\xi_{M,n}\!\\ \end{bmatrix}\!}_{\triangleq\,\bm{\xi}_{n}\,\in\,\mathbb{R}^{M\!\times\!1}}\!\!\!\!\!\!\!\in\!\mathbb{R}^{M\times 1}\!.$
In the latter representation of the observed data vector $\bm{y}_{n}\in\mathbb{R}^{M\times 1}$, we have implicitly defined an effective channel matrix and $\bm{G}$ $\in\mathbb{R}^{M\times 4}$, which is constructed from range measurements and anchor positions; the vector $\bm{x}_{n}\in\mathbb{R}^{4\times 1}$ which contains the unknown coordinates of the $n$-sensor, as well as its distance to the origin; and a vector $\bm{\xi}_{n}\in\mathbb{R}^{M\times 1}$, which gathers the composite noise quantities defined in equation (16).
As hinted in the introduction, the linear system in equation (LABEL:eq:linear_sys) can be leveraged for the estimation of the unknown sensor coordinate vector $\bm{s}_{n}$ and sensor position norm $||\bm{s}||_{2}^{2}$ in $\bm{x}_{n}$, from which equation (11) can be invoked for the translation and rotation extraction via Procrustes analysis or other classical algorithms [9]. We shall return to this idea in Section III.
@@ II-C Sensor Doppler Measurement Model
which can also be reformulated in terms of a composite Doppler noise $\zeta_{n}\in\mathbb{R}$, given by
@@ III Proposed Rigid Body Localization
In this section, we propose a low-complexity position and transformation estimator for RBL, in light of the system model derived in Section II and by leveraging the GaBP message passing framework. Specifically, an initial GaBP is derived to solve the position-explicit system in equation (LABEL:eq:linear_sys) to obtain sensor coordinates. Then, in possession of the initial position estimate, a second GaBP is derived on the rigid body transformation parameter-explicit system in equation (28), to obtain the final estimate of the 3D rotation angles $\bm{\theta}$ and translation vector $\bm{t}$.
@@ III-A Stationary Transformation Parameter-based System Model
In this section, we build upon equation (LABEL:eq:linear_sys) towards a reformulation that expresses system variables directly in terms of the RBL transformation parameters, i.e., the 3D rotation angles $\bm{\theta}\triangleq[\theta_{x},\theta_{y},\theta_{z}]^{\mathsf{T}}\in\mathbb{R}^{3\times 1}$ and translation vector $\bm{t}$ [12]. To that end, we first apply a small-angle approximation [^4] For practical rigid body tracking applications, subsequent transformation estimations can be assumed to be performed within a sufficiently short time period, such that the change in rotation angle is small. [10] onto the rotation matrix of equation (10), obtained by leveraging $\cos\theta\approx 1$ and $\sin\theta\approx\theta$, which yields
@@
$+\underbrace{\begin{bmatrix}0&1&0&-1&0&0&0&0&0\\ 0&0&-1&0&0&0&1&0&0\\ 0&0&0&0&0&1&0&-1&0\\ \end{bmatrix}^{\!\mathsf{T}}\!\!\!}_{\triangleq\,\bm{L}\,\in\,\mathbb{R}^{9\times 3}}\cdot\!\begin{bmatrix}\theta_{x}\\ \theta_{y}\\ \theta_{z}\end{bmatrix}\!\!.$
@@
$\xi_{n}\!=$ $\tilde{d}_{m,n}^{2}\!-\!\left\|\bm{a}_{m}\right\|^{2}_{2}\!-\!\left\|\bm{s}_{n}\right\|^{2}_{2}\!+\!2\!\left[\bm{c}_{n}^{\mathsf{T}}\otimes\bm{a}_{m}^{\mathsf{T}}\right]\!\bm{\gamma}\!+\!2\!\left[\bm{c}_{n}^{\mathsf{T}}\otimes\bm{a}_{m}^{\mathsf{T}}\right]\!\bm{L}\bm{\theta}$ (27)
$+2\bm{a}_{m}^{\mathsf{T}}\bm{t},\in\mathbb{R},\!\!$
@@
and
$\bm{H}_{\theta}\!=\!\!\left[\begin{array}[]{c}\!\!\!\!\!-2\!\left[\bm{c}_{i}^{\mathsf{T}}\otimes\bm{a}_{1}^{\mathsf{T}}\right]\!\bm{L}\!\!\!\\ \vdots\\[2.15277pt] \!\!\!\!-2\!\left[\bm{c}_{i}^{\mathsf{T}}\otimes\bm{a}_{M}^{\mathsf{T}}\right]\!\bm{L}\!\!\!\end{array}\right]\!\!\in\!\mathbb{R}^{M\times 3}\!,~{}\bm{H}_{t}\!=\!\!\left[\begin{array}[]{c}\!\!\!\!-2\bm{a}_{1}^{\mathsf{T}}\!\!\!\!\\ \vdots\\ \!\!\!\!-2\bm{a}_{M}^{\mathsf{T}}\!\!\!\end{array}\right]\!\!\in\!\mathbb{R}^{M\times 3}\!,$ (28i)
where $\bm{z}_{n}\in\mathbb{R}^{M\times 1}$ is the effective observed data vector, and $\bm{H}_{\theta}\in\mathbb{R}^{M\times 3}$ and $\bm{H}_{t}\in\mathbb{R}^{M\times 3}$ are respectively the effective channel matrices for the unknown rotation and translation variables, and $\bm{\xi}_{n}\in\mathbb{R}^{M\times 1}$ is the vector of composite noise variables from equation (16).
@@ III-B Linear GaBP for Sensor Position Estimation
In what follows, we will derive the step-by-step GaBP message-passing rules for RBL based on the linear model of equation (LABEL:eq:linear_sys). Since the derivation is identical for each $n$-th sensor node, we shall for the sake of convenience temporarily drop the subscript n in the quantities $\bm{y}_{n}$, $\bm{x}_{n}$ and $\bm{\xi}_{n}$ in equation (LABEL:eq:linear_sys), such that we can define the elements of each of these vectors respectively as $\bm{y}_{n}\to\bm{y}=\left[y_{1},\ldots,y_{m},\ldots,y_{M}\right]^{\mathsf{T}},\in\mathbb{R}^{M\times 1}$, $\bm{x}_{n}\to\bm{x}=\left[x_{1},\ldots,x_{k},\ldots,x_{K+1}\right]^{\mathsf{T}},\in\mathbb{R}^{K+1\times 1}$, with $k=\{1,\ldots,K\!+\!1\}$ where $K$ is the dimension of the space, and $\bm{\xi}_{n}\to\bm{\xi}=\left[\xi_{1},\ldots,\xi_{m},\ldots,\xi_{M}\right]^{\mathsf{T}},\in\mathbb{R}^{M\times 1}$. In addition, let $g_{m,k}$ denote the entries of the matrix $\bm{G}$ $\in\mathbb{R}^{M\times K+1}$.
@@
where $\tilde{y}_{m,k}^{[j]}\in\mathbb{R}$ is the soft-IC symbol corresponding to the $m$-th element of the observed data vector and the $k$-th element of the position vector, and $g_{m,k}\in\mathbb{R}$ is the $(m,k)$-th element of the effective channel matrix $\bm{G}$ in equation (16).
@@
The extrinsic mean and variance are subsequently denoised with a classic zero-mean Gaussian denoiser [^5] If an alterative prior distribution of the position variables are assumed, i.e., uniform distribution, a different Bayes-optimal denoiser can be utilized., such that the denoised mean and variance are given by
$\check{x}_{m,k}^{[j]}=\frac{\phi_{x}\bar{x}_{m,k}^{[j]}}{\phi_{x}+\bar{v}_{m,k}^{[j]}}\in\mathbb{R}~{}\,\text{and}\,~{}\check{\psi}_{m,k}^{[j]}=\frac{\phi_{x}\bar{v}_{m,k}^{[j]}}{\phi_{x}+\bar{v}_{m,k}^{[j]}}\in\mathbb{R},\!\!\!$ (35)
where $\phi_{x}\in\mathbb{R}$ is the variance of the prior distribution of the position variable.
Finally, the $(j\!+\!1)$-th soft-replica is obtained via the following damped update to prevent error floors caused by early erroneous convergence to a local optima
where $\rho\in[0,1]$ is a selected damping parameter.
@@
Algorithm 1 summarizes the proposed method for the sensor position estimation without rigid body conformation, based on the range information between the sensors and anchors, ultimately yielding $(K+1)$ elements in $\bm{x}_{n}$ composed of $K$ coordinate estimates of $\mathbf{s}_{n}$ and its absolute norm $||\bm{s}_{n}||_{2}^{2}$.

### v3

# Please find the official IEEE published version of this article on IEEE Xplore [here] and cite as: N Führling et al., “6D Rigid Body Localization and Velocity Estimation via Gaussian Belief Propagation,” in IEEE Transactions on Signal Processing, vol. 73, pp. 3902-3917, Sept. 2025

@@ I Introduction
Several effective strategies exist to estimate the location and attitude of objects, including computer vision-based techniques, which focus on feature extraction and posture estimation [17, 18], usually based on image/video signals; and inertial measurement unit (IMU)-based techniques, which leverage information from accelerometers, gyroscopes, and magnetometers [19, 20]. A problem of computer-vision approaches is, however, that they typically require high volumes of data and rely on high-complexity methods, which limits their application. In turn, IMU-based methods are often unreliable or inaccurate, requiring frequent sensor re-calibrations, not to mention the aid of external radio technology, which to some extent defeats the self-reliant idea behind the approach.
In contrast to the latter, the RBL concept considered here [10, 11, 12, 13, 14, 15] aims to generalize the traditional localization paradigm, and exploits range measurements between sensors on the rigid body and a set of anchor sensors at known positions [21, 22], to estimate the location, orientation (and possibly the shape) of rigid bodies. Earlier state-of-the-art (SotA) contributions to such RBL approaches are, however, based on either algebraic methods leveraging multidimensional scaling (MDS) [23] or semidefinite relaxation (SDR) [24, 25], both of which have a cubic computational complexity at least. And although least squares methods [26, 27, 28] have been used to lower the complexity of the RBL problem, these techniques have in common the fact that they only take into account stationary scenarios, in the sense that they seek to estimate only the translation and the orientation ($i.e.$, the pose) of the rigid body, respectively defined as the distance and the rotation matrix of the rigid body relative to a given reference.
@@
The capitalizing feature of the method in [12] to handle both the stationary and moving RBL problems similarly is thanks to the linearized error model described in [31] and [32]. However, the least-squares formulation and two stage approach adopted can limit performance [33] compared to message-passing methods [34], which are well known to achieve low complexity while maintaining a good performance especially in problems with coupled variables, which can be solved not only via alternating methods [35], but also jointly via bilinear techniques [36, 37, 38, 39].
Among existing localization approaches based on message passing techniques, it should be noted that, to the best of the authors’ knowledge, the only method specifically designed for RBL is the work introduced in [2]. Nevertheless, several other message passing-based localization techniques, though not specifically designed for RBL, are worth presenting. For example, the method presented in [40] employs Gaussian belief propagation (GaBP) for simultaneous localization and synchronization based on time of arrival (ToA) measurements. Additionally, in [41], message passing algorithms are exploited to solve a probabilistic inference problem in received signal strength indicator (RSSI)-based cooperative localization scenarios involving an unknown path loss exponent. Moreover, the approach described in [42] leverages GaBP for distributed multiple-input multiple-output (MIMO) radar localization, resulting in low computational complexity and reduced communication cost.
Motivated by the aforementioned studies, and the identification of the lack of message passing-based RBL techniques – a series of novel algorithms are proposed in this article to address both stationary and moving RBL problems. First, two methods for the location and velocity estimation of sensors in three-dimensional (3D) space are produced. In particular, the techniques, labeled Algorithms 1 and 3, respectively, are designed for the estimation of sensor positions from range measurements, as well as sensor velocities from Doppler measurements. Then, two bivariate linear GaBP algorithms are offered for the stationary and moving RBL problems. Those contributions, dubbed Algorithms 2 and 4, respectively, are capable of estimating the 3D rotation angles and translation distances, and respectively the 3D angular and translational velocities, utilizing range/Doppler measurements and the estimated sensor position and velocities from Algorithms 1 and 2 by leveraging bilinear GaBP over a small-angle modification of the linearized system models of [32] and [31], with the rigid body conformation constraint directly incorporated. The bivariate methods in Algorithms 4 and 3, respectively, are shown to outperform the SotA two-stage RBL methods in terms of parameter estimation performance, while retaining low computational complexity.
@@ II-A Rigid Body System Model
Figure 2: An illustration of a moving rigid body, whose location has a 3D rotation $\bm{Q}$ and a translation $\bm{t}$ from the reference frame33footnotemark: 3, as determined by equation (11). The velocity $\bm{\dot{s}}_{n}$ of the $n$-th sensor as a function of the angular velocity $\bm{\omega}$ and translational velocity $\bm{\dot{t}}$, as determined by equation (12), is also illustrated.
@@
The latter system model for stationary rigid bodies can be extended to moving rigid bodies [^4] It should be noted that the employed model is purely kinematic in nature and does not incorporate additional physical aspects such as the forces acting upon the rigid body or its mass. Furthermore, due to the adopted kinematic formulation, all parameters are considered in a decoupled manner. This simplification can be revisited in a future work, where a coupled model may be adopted and analyzed. by introducing the angular velocity vector $\bm{\omega}\triangleq\left[\omega_{1},\omega_{2},\omega_{3}\right]^{\mathsf{T}}\in\mathbb{R}^{3\times 1}$, and the translational velocity vector $\bm{\dot{t}}\triangleq[\dot{t}_{x},\dot{t}_{y},\dot{t}_{z}]^{\mathsf{T}}\in\mathbb{R}^{3\times 1}$, such that the velocity of the $n$-th sensor can be expressed as
@@
where $\cdot$$^{\times}$ is the cross product operator for matrices [11], which maps the angular velocity vector $\bm{\omega}$ to a skew-symmetric matrix, given by
$\bm{\omega}]^{\times}=\left[\begin{array}[]{ccc}0&-\omega_{3}&\omega_{2}\\ \omega_{3}&0&-\omega_{1}\\ -\omega_{2}&\omega_{1}&0\end{array}\right]\in\mathbb{R}^{3\times 3}.$ (13)
@@ II-B Sensor Range Measurement Model
44footnotetext: The reference frame is generally not at the origin, but since a rigid body rotation is only relative to its previous orientation, setting the reference frame to the origin, $i.e.$, making $\bm{Q}^{(0)}\triangleq\mathbf{I}_{3\times 3}$, can be done without loss of generality (w.l.g.).
@@
where the second-order noise term $w_{m,n}^{2}$ is neglected [^5] The second-order noise terms, often neglected in linear approximations, can introduce non-negligible biases in nonlinear estimation problems. Although their influence is typically minor in large scale systems, such as the scenario considered in this work, it may become significant primarily in small scale systems involving nonlinear estimation under high noise conditions..
@@
[…]
In the latter representation of the observed data vector $\bm{y}_{n}\in\mathbb{R}^{M\times 1}$, the effective channel matrix $\bm{G}$ $\in\mathbb{R}^{M\times 4}$ has been implicitly defined to be constructed from range measurements and anchor positions $\bm{a}_{m}$; the vector $\bm{x}_{n}\in\mathbb{R}^{4\times 1}$ which contains the unknown coordinates of the $n$-sensor, as well as its distance to the origin; and a vector $\bm{\xi}_{n}\in\mathbb{R}^{M\times 1}$, which gathers the composite noise quantities defined in equation (16).
As hinted in the introduction, the linear system in equation (LABEL:eq:linear_sys) can be leveraged for the estimation of the unknown sensor coordinate vector $\bm{s}_{n}$ and sensor position norm $\bm{s}||_{2}^{2}$ in $\bm{x}_{n}$, from which equation (11) can be invoked for the translation and rotation extraction via Procrustes analysis or other classical algorithms [10]. We shall return to this idea in Section III.
@@ II-C Sensor Doppler Measurement Model
which can also be reformulated in terms of a composite Doppler noise $\zeta_{{\color[rgb]{0,0,0}m,}n}\in\mathbb{R}$, given by
@@ III Proposed Rigid Body Localization
In this section, we propose a low-complexity position and transformation estimator for RBL, in light of the system model derived in Section II and by leveraging the GaBP message passing framework. Specifically, an initial GaBP is derived to solve the position-explicit system in equation (LABEL:eq:linear_sys) to obtain jointly the entire 3D sensor coordinates. Then, in possession of the initial position estimate, a second GaBP is derived on the rigid body transformation parameter-explicit system in equation (28), to obtain the final estimate of the 3D rotation angles $\bm{\theta}$ and translation vector $\bm{t}$.
@@ III-A Stationary Transformation Parameter-based System Model
In this section, we build upon equation (LABEL:eq:linear_sys) towards a reformulation that expresses system variables directly in terms of the RBL transformation parameters, i.e., the 3D rotation angles $\bm{\theta}\triangleq[\theta_{x},\theta_{y},\theta_{z}]^{\mathsf{T}}\in\mathbb{R}^{3\times 1}$ and translation vector $\bm{t}$ [13]. To that end, we first apply a small-angle approximation [^6] For practical rigid body tracking applications, subsequent transformation estimations can be assumed to be performed within a sufficiently short time interval such that the change in rotation angle remains small. Although the approximation remains valid for rotation angles up to approximately twenty degrees, convergence of the algorithm can still be achieved for larger angles. However, in such cases, the estimation accuracy may degrade. [11] onto the rotation matrix of equation (10), obtained by leveraging $\cos\theta\approx 1$ and $\sin\theta\approx\theta$, which yields
@@
$\hskip-206.66602pt+\underbrace{\begin{bmatrix}0&1&0&-1&0&0&0&0&0\\ 0&0&-1&0&0&0&1&0&0\\ 0&0&0&0&0&1&0&-1&0\\ \end{bmatrix}^{\!\mathsf{T}}\!\!\!}_{\triangleq\,\bm{L}\,\in\,\mathbb{R}^{9\times 3}}\cdot\!\begin{bmatrix}\theta_{x}\\ \theta_{y}\\ \theta_{z}\end{bmatrix}\!\!$.
@@
$\xi_{n}\!=$ $\hskip-17.22217pt\tilde{d}_{m,n}^{2}\!-\!\left\|\bm{a}_{m}\right\|^{2}_{2}\!-\!\left\|\bm{s}_{n}\right\|^{2}_{2}\!+\!2\!\left[\bm{c}_{n}^{\mathsf{T}}\otimes\bm{a}_{m}^{\mathsf{T}}\right]\!\bm{\gamma}\!+\!2\!\left[\bm{c}_{n}^{\mathsf{T}}\otimes\bm{a}_{m}^{\mathsf{T}}\right]\!\bm{L}\bm{\theta}$ (27)
$\hskip-17.22217pt+2\bm{a}_{m}^{\mathsf{T}}\bm{t},\in\mathbb{R},\!\!$
@@
where $\bm{z}_{n}\in\mathbb{R}^{M\times 1}$ is the effective observed data vector, $\bm{\xi}_{n}\in\mathbb{R}^{M\times 1}$ is the vector of composite noise variables from equation (16), and the two effective channel matrices for rotation and translation respectively defined as
$\color[rgb]{0,0,0}\bm{H}_{\theta}\!=\!\!\left[\begin{array}[]{c}\!\!\!\!\!-2\!\left[\bm{c}_{i}^{\mathsf{T}}\otimes\bm{a}_{1}^{\mathsf{T}}\right]\!\bm{L}\!\!\!\\ \vdots\\[2.15277pt] \!\!\!\!-2\!\left[\bm{c}_{i}^{\mathsf{T}}\otimes\bm{a}_{M}^{\mathsf{T}}\right]\!\bm{L}\!\!\!\end{array}\right]\!\!\in\!\mathbb{R}^{M\times 3}\!,~\bm{H}_{t}\!=\!\!\left[\begin{array}[]{c}\!\!\!\!-2\bm{a}_{1}^{\mathsf{T}}\!\!\!\!\\ \vdots\\ \!\!\!\!-2\bm{a}_{M}^{\mathsf{T}}\!\!\!\end{array}\right]\!\!\in\!\mathbb{R}^{M\times 3}.$ (28i)
@@ III-B Linear GaBP for Sensor Position Estimation
In what follows, we will derive the step-by-step GaBP message-passing rules for RBL based on the linear model of equation (LABEL:eq:linear_sys). Since the derivation is identical for each $n$-th sensor node, we shall for the sake of convenience temporarily drop the subscript n in the quantities $\bm{y}_{n}$, $\bm{x}_{n}$ and $\bm{\xi}_{n}$ in equation (LABEL:eq:linear_sys), such that we can define the elements of each of these vectors respectively as $\bm{y}_{n}\to\bm{y}=\left[y_{1},\ldots,y_{m},\ldots,y_{M}\right]^{\mathsf{T}},\in\mathbb{R}^{M\times 1}$, $\bm{x}_{n}\to\bm{x}=\left[x_{1},\ldots,x_{k},\ldots,x_{K+1}\right]^{\mathsf{T}},\in\mathbb{R}^{K+1\times 1}$, with $k=\{1,\ldots,K\!+\!1\}$ where $K$ is the dimension of the space, and $\bm{\xi}_{n}\to\bm{\xi}=\left[\xi_{1},\ldots,\xi_{m},\ldots,\xi_{M}\right]^{\mathsf{T}},\in\mathbb{R}^{M\times 1}$. In addition, let $g_{m,k}$ denote the entries of the matrix $\bm{G}$.
The proposed GaBP algorithm operates on a linear system of the form $\bm{y}=\bm{G}\bm{x}+\bm{\xi}$, where $\bm{x}$ denotes the collection of unknown variables, such as sensor positions, velocities, or rigid body parameters as discussed in the following sections. The matrix $\bm{G}$ represents the system matrix determined by the geometric configuration of anchors and sensors.
The associated factor graph is bipartite, comprising variable nodes corresponding to each element of $\bm{x}$ and factor nodes for each observation element of $\bm{y}$, with connections defined by the nonzero entries in $\bm{G}$, as illustrated in Figure 3. As evident from the system formulation in the previous section, this graph structure is sparse in practical deployments and the resulting system is well conditioned [^7] In all scenarios considered in the following sections, the system matrices are well conditioned given appropriate physical layouts of sensors and anchors. This characteristic is known to influence the convergence of GaBP, subject to conditions such as walk summability or diagonal dominance of the information matrix. As detailed in Section V, where a comprehensive performance analysis of the proposed GaBP algorithms is provided, empirical evaluations have demonstrated consistent convergence across all tested configurations. under suitable arrangements.
@@
Figure 3: Factor graph structure of the linear system in equation (LABEL:eq:linear_sys), where the variable nodes (circles) represent the unknown variables $x_{k}$ and the factor nodes (squares) represent the observations $y_{m}$. Edges are defined by the connecting entries of the effective channel matrix $\bm{G}$.
@@
where $\tilde{y}_{m,k}^{[j]}\in\mathbb{R}$ is the soft-IC symbol corresponding to the $m$-th element of the observed data vector and the $k$-th element of the position vector, and $g_{m,k}\in\mathbb{R}$ is the $(m,k)$-th element of the effective channel matrix $\bm{G}$ in equation (LABEL:eq:linear_sys).
@@
The extrinsic mean $\bar{x}_{m,k}^{[j]}$ and variance $\bar{v}_{m,k}^{[j]}$ are subsequently combined with the prior distribution [^8] The statistics of the prior distribution of the RBL parameters are assumed to be known, which is a common assumption in the literature [12, 30], but can also be learned from the data within the iterations of the GaBP via methods such as expectation-maximization approaches [45] if they are assumed to be not available a priori, where the latter is out of scope of this article. of the RBL position variables to obtain the denoised posterior estimates.
In accordance with the GaBP framework, the Bayes-optimal denoiser is derived from the product of the extrinsic PDF of the position variable and the prior distribution, which are both assumed to be Gaussian [^9] If an alternative prior distribution of the position variables are assumed, i.e., uniform distribution, a different Bayes-optimal denoiser can be utilized., such that the posterior distribution is Gaussian with mean $\check{x}_{m,k}^{[j]}$ and variance $\check{\psi}_{m,k}^{[j]}$ given by
$\check{x}_{m,k}^{[j]}=\frac{\bar{v}_{m,k}^{[j]}\mu_{x}+\phi_{x}\bar{x}_{m,k}^{[j]}}{\phi_{x}+\bar{v}_{m,k}^{[j]}}\text{ and }\check{\psi}_{m,k}^{[j]}=\frac{\phi_{x}\bar{v}_{m,k}^{[j]}}{\phi_{x}+\bar{v}_{m,k}^{[j]}}$ (35)
@@
with $\mu_{x}\in\mathbb{R}$ and $\phi_{x}\in\mathbb{R}$ being the mean and variance of the prior distribution of the position variable, where since it is assumed $\mu_{x}=0$, the final denoised posterior estimates are simplified to
$\check{x}_{m,k}^{[j]}=\frac{\phi_{x}\bar{x}_{m,k}^{[j]}}{\phi_{x}+\bar{v}_{m,k}^{[j]}}\in\mathbb{R}~\,\text{and}\,~\check{\psi}_{m,k}^{[j]}=\frac{\phi_{x}\bar{v}_{m,k}^{[j]}}{\phi_{x}+\bar{v}_{m,k}^{[j]}}\in\mathbb{R}.\!\!\!$ (36)
Finally, the (j+1)(j\!+\!1)-th soft-replica is obtained via the following damped update to prevent error floors caused by early erroneous convergence to a local optima
where $\rho\in[0,1]$ is a selected damping parameter.
@@
Algorithm 1 summarizes the proposed method for the sensor position estimation without rigid body conformation, based on the range information between the sensors and anchors, ultimately yielding $(K+1)$ elements in $\bm{x}_{n}$ composed of $K$ coordinate estimates of $\mathbf{s}_{n}$ and its absolute norm $\bm{s}_{n}||_{2}^{2}$.